15 - 3 × 6 ÷ 1 + 2
Answer by lenny460(1073)

15 - 3 × 6 ÷ 1 + 2
We use PEMDAS or BODMAS to solve the above.
-3 x 6 = -18
15 - 18 / 1 + 2
-18 / 1 = -18
15 - 18 + 2
15 - 18 = -3
-3 + 2 = -1

The Answer:
-1